The following selected data were taken from the accounting records of Colorado Enterprises:

Month Machine Hours Manufacturing Overhead

May 49,500 $ 924,000

June 61,400 1,144,000

July 75,000 1,348,700

August 55,500 987,000

Manufacturing overhead consists of three different costs; (1) machine supplies (variable), (2) property taxes (fixed), and (3) plant maintenance (semivariable). July's overhead costs were $174,000 for machine supplies, $24,700 for property taxes, and $1,150,000 for plant maintenance.

Required: Determine the machine supplies and property taxes for May. By using the high-low method, analyze Colorado’s plant maintenance cost and calculate the monthly fixed portion and the variable cost per machine hour. Assume that present cost behavior patterns continue into future months. Estimate the total amount of manufacturing overhead the company can expect in September if 57,400 machine hours are worked.
Answer #1

a) MACHINE SUPPLIES AND PROPERTY TAXES FOR THE MONTH OF MAY :

Machine supplies(May)

= (machine supplies overheads for July ÷ machine hrs of July)× machine hrs of May

= ($174000÷75000)×49500

= $2.32 × 49500

= $114840.

Property taxes for July is given as $24700 are fixed cost and remains same every month.

Therefore, property taxes for month of may = $24700

b) MONTHLY FIXED PORTION AND VARIABLE COST PER MACHINE HOUR FOR PLANT MAINTAINANCE :

Plant Maintainance in May

= Mfg. Overhead of May - property taxes of may - machine supplies of may

= $924000 - $24700 - $114840

= $784460.

VARIABLE plant maintainance

= ( Difference in plant maintaince cost of july and may ÷ Difference in machine hrs of july and may )

= ( $1150000 - $784460 ) ÷ ( 75000 - 49500 )

= $365540 ÷ 25500

= $14.33 per machine hour.

FIXED portion of plant maintainance

= Total plant maintainance for 75000 hrs - Variable plant maintainance for 75000 hrs.

= $1150000 - ( 75000 hrs. × $14.33 )

= $1150000 - $1074750

= $75250.

c) ESTIMATION OF MANFACTURING OVERHEAD FOR THE MONTH OF SEPTEMBER IF 56000 MACHINE HOURS ARE WORKED :

Machine supplies for 56000 hrs = ($174000÷75000)×56000 = $2.32×56000 = $129920.

Property taxes ( fixed ) = $24700.

Plant Maintainance ( variable ) = $14.33 × 56000 = $802480.

Plant Maintainance ( fixed ) = $75250

Total Manufacturing overhead

= $129920 + $24700 + $802480 + $75250

= $1032350.

_____×_____

ALL THE BEST

Keep Learning